A Restaurant manager decides to lower the price of a featured sandwich from Rs. 500


to Rs. 400, and she finds that sales during the week increase from 440 to 680


sandwiches. Is demand elastic or Inelastic?


1
Expert's answer
2021-05-04T07:20:47-0400

Initial Price =500

New Price =400

Initial Quantity =440

New Quantity =680


"PED =\frac {\Delta Q} {\Delta P} ×(\frac {P_2+P_2}{Q_1+Q_2})"


"PED =\frac {680-440} {400-500} ×(\frac {500+400}{440+680})"


"PED=\frac {240}{-100}×\frac {900}{1120}"


"PED=-2.4\u00d70.8036"


"PED=-1.93"


Hence the demand for sandwiches is elastic.